### v1

# Quantum Oppenheimer-Snyder models in loop quantum cosmology with Lorentz term

Abstract
A novel quantum black hole model is derived by incorporating the Lorentzian term within the loop quantum cosmology framework of the quantum Oppenheimer-Snyder (qOS) model. This model features a quantum-corrected metric tensor, representing a deformation of the classical Schwarzschild solution. Investigations into the quasi-normal modes reveal that these quantum-corrected black holes exhibit stability against scalar perturbations. Notably, the exponential decay rate within the Lorentzian qOS model demonstrates a significant reduction compared to both the earlier qOS model devoid of this term and the standard Schwarzschild black hole. The higher overtones of the Lorentzian qOS model also differ significantly from those of the earlier qOS model and the standard Schwarzschild black hole, indicating that the near-horizon geometry is substantially modified. The thermodynamic properties with both positive and negative cosmological constant are considered. For the anti-de Sitter case, our analysis reveals that for small black hole masses, the temperature within this loop quantum gravity framework decreases as the mass diminishes, contrasting with classical black hole behavior. Furthermore, a logarithmic term emerges as the leading-order correction to the Bekenstein-Hawking entropy. Additionally, LQG corrections induce an extra phase transition in the black hole’s heat capacity at smaller radius. While for the de Sitter case, the temperature increases as the mass decreases for small black holes, again differing from classical expectations. Similarly to the anti-de Sitter case, LQG corrections in the de Sitter case also lead to an extra phase transition in the heat capacity at small radius.
I Introduction
General Relativity (GR) encounters fundamental limitations when addressing spacetime singularities, particularly where curvature approaches Planck-scale magnitudes. This breakdown necessitates a quantum gravity framework capable of unifying quantum mechanics with GR to resolve these singularities. Among promising candidates, Loop Quantum Gravity (LQG) stands out due to its background-independent, nonperturbative nature and mathematical rigor [1, 2, 3, 4].
Recent advancements in LQG have yielded significant breakthroughs, particularly in resolving the Big Bang singularity and elucidating Hawking-Bekenstein black hole entropy [5, 6, 7]. The application of LQG principles to homogeneous, isotropic universes has established Loop Quantum Cosmology (LQC), whose hallmark achievement is replacing the classical Big Bang singularity with a quantum bounce mechanism [8, 9, 10, 11, 12, 13, 14, 15].
In classical gravitational theory, the Oppenheimer-Snyder (OS) model [16] provides a foundational framework for studying gravitational collapse through pressureless, homogeneous dust dynamics governed by Friedmann equations. While valuable, this model inherits GR’s singularity problem. Recent work has addressed this limitation by applying LQC corrections to the OS framework, resulting in the quantum Oppenheimer-Snyder (qOS) model [17], which has subsequently undergone extensive theoretical investigation [18, 19, 20].
Within the current qOS framework, the interior dynamics of black holes are predominantly analyzed using the simplest Loop Quantum Cosmology (LQC) approach, specifically the Ashtekar-Pawlowski-Singh (APS) model [11]. It is noteworthy that the Hamiltonian constraint in Loop Quantum Gravity (LQG) comprises both Euclidean and Lorentzian components. Classically, these terms exhibit proportionality, yet the APS model employs a pre-quantization reduction that retains only the Euclidean term [9]. This simplification yields a symmetric bounce behavior during both pre- and post-bounce phases [11].
Emerging research indicates that this conventional treatment fails to fully capture LQG’s complete feature set. Crucially, full theory quantization demonstrates distinct behaviors between Euclidean and Lorentzian terms [21, 22, 23, 24]. Through Thiemann’s regularization technique, a refined LQC model incorporating the Lorentzian term has been developed. This advancement introduces significant modifications to the Friedmann equation [25, 26] manifesting an emergent de-Sitter epoch and asymmetric bounce dynamics. This raises a pivotal question: Would substituting the APS model with this Lorentzian-enhanced LQC preserve the qOS framework’s core conclusions? Our investigation addresses this through systematic derivation and analysis the physical properties of the modified qOS-LQC model.
One important physical property is that the black hole perturbation dynamics: When subjected to external perturbations, black holes exhibit a distinctive gravitational wave emission pattern that unfolds in three sequential stages. The initial burst phase, whose intensity and duration depend on the perturbation conditions, is followed by the quasinormal mode (QNM) ringdown - a hallmark feature where spacetime oscillations decay with characteristic complex frequencies. These QNM signatures, recently confirmed through groundbreaking gravitational wave detections [27, 28] have become powerful tools for probing the fundamental geometry of black holes. The ringdown profile’s precision is governed by both the fundamental mode ($n=0$) and its higher-order overtones ($n>0$) [29, 30, 31] with each overtone contributing unique information about the black hole’s structure. The final phase, known as late-time tail evolution, displays either power-law or exponential decay patterns that encode crucial information about the black hole’s properties [32], particularly regarding the nature of its Cauchy horizon [33, 34]. Consequently, the QNMs and the late-time tail, are applied to test the stability of black holes and encode important information about the black holes. To this end, we employ the finite element method and WKB approximation to investigate the QNMs, overtone, and late-time tails of scalar perturbations in quantum black holes.
Recently, black hole thermodynamics has attracted great interest. The Hawking radiation paradigm [35, 36], established black holes as thermodynamic entities possessing temperature and entropy. The four laws of black hole thermodynamics [37], parallel classical thermodynamic principles, with Bekenstein-Hawking entropy bridging classical and quantum gravitational theories [38].
Concurrently, cosmological observations of universal acceleration necessitate dark energy models, where the cosmological constant $\Lambda$ remains the predominant explanation [39, 40]. In particular, the recent cosmological observations of the Universe are consistent with a de Sitter (dS) spacetime, characterized by a positive cosmological constant ($\Lambda>0$). Meanwhile, the success of the AdS/CFT correspondence has highlighted the theoretical importance of anti-de Sitter (AdS) spacetime, characterized by a negative cosmological constant ($\Lambda<0$). The AdS black holes exhibit unique thermodynamic behaviors, such as phase transitions and critical phenomena [41, 42, 43], which are not observed in asymptotically flat spacetimes. One of the most interesting black hole phase transitions is the Hawking-Page phase transition [44]. So, in this work, we focus on investigating the thermodynamic properties of black holes in both de Sitter and anti-de Sitter spacetimes.
This paper is organized as follows. Firstly, we derive the new model by combining the modified Friedmann equation with the qOS model, and analyze its physical properties in Sect. II. Then, we investigate the QNMs, the late-time tail behavior, and the behavior of higher overtones in Sect. III. In the Sect. IV, we examine the thermodynamic properties of the model. Finally, we summarize and discuss our main results in Sect. V. Throughout this work, we use geometrical units in which $G=c=\hbar=1$.
II The qOS model in LQC with Lorentz term
The qOS model’s spacetime comprises the pressureless dust ball and the vacuum region outside the dustball, the metric tensor inside of the dustball can be described by the FRW metric as
$ds_{in}^{2}=-d\tau^{2}+a(\tau)^{2}(d\tilde{r}^{2}+d\Omega^{2}),$ (1)
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\ d\phi^{2}$, and $a(\tau)$ is the scale factor of the universe. The original qOS model uses the Ashtekar-Pawlowski-Singh (APS) model [11] deciping. However, the APS version of LQC model only consider the Euclidean term in Hamiltonian constraint and stand for a simplified model. The recent advances of LQC found that the Lorentz term should also need to be considered and will bring the significant changes for the cosmic evolution such as the de-Sitter epoch will emerged. The dynamics of the $a(\tau)$ of LQC with Lorentz term governed by the modified Friedmann equation [25, 26]
$H^{2}=(\frac{\dot{a}}{a})^{2}=\frac{1}{\gamma^{2}\Delta}f(\rho)(1-f(\rho))\left(1-\frac{\rho}{\rho_{c}}\right),$ (2)
where
$f(\rho)=\frac{1\pm\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi\tilde{r}^{3}a^{3}}.$
Here $\Delta=2\pi\sqrt{3}\gamma G\hbar$, which is the smallest non-vanishing area eigenvalue from the full theory [1, 2]. The $\gamma$ is the Barbero-Immirzi parameter. Using the black hole thermodynamics in LQG, so we could set the value being 0.2375. ${\rho_{c}}=\frac{3}{32\pi G\Delta\gamma^{2}(1+\gamma^{2})}$ is the critical energy density of this model. In the classical limit $\frac{\rho}{\rho_{c}}\ll 1$, the (2) reduce to the standard Friedmann equation. When $\rho=\rho_{c}$, the quantum bounce occurs and divides the evolution of the universe into asymmetric periods.
We should note that the modified Friedmann equation has two branches. When $f(\rho)=\frac{1+\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$, the corresponding modified Friedmann equation describes the pre-bounce epoch. Although the classical model does not contain a cosmological constant, the quantum geometric effects are equivalent to produce a positive effective cosmological constant of Planckian order, resulting in a de Sitter epoch. Another branch is given by $f(\rho)=\frac{1-\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$, which corresponds to the post-bounce epoch. This epoch is characterized by a classical universe without a cosmological constant, corresponding to the universe where we live. So in this work, we mainly focus on the post-bounce sector.
Outside the dust ball’s region, we assume that the spacetime is spherically symmetric and static, so the metric reads
$ds_{out}^{2}=-F(r)dt^{2}+G(r)^{-1}dr^{2}+d\Omega^{2}.$ (3)
To determine the unknown functions $F(r)$ and $G(r)$, we need to consider the Darmois-Israel junction condition [45, 46]. Note that the coordinates $(\tau,\tilde{r},\theta,\phi)$ describe the interior of a dust ball with a radius $0<\tilde{r}<\tilde{r}_{0}$, while the exterior is described by the coordinates $(t,r,\theta,\phi)$, with $r$ being the radial coordinate in the exterior region. To construct the complete spacetime, we need to glue the interior region described by the dust ball to the innermost region of a spherically symmetric and static exterior spacetime. In the interior region, the geodesic of a freely falling particle can be described by $\tau\mapsto(\tau,\tilde{r}_{0},\theta,\phi)$, while in the exterior region, it is described by $\tau\mapsto(t(\tau),r(\tau),\theta,\phi)$, so the single surfaces are glued by the identification $(\tau,\tilde{r}_{0},\theta,\phi)\sim(t,r,\theta,\phi)$. From the geodesic, we can derive:
$-\bigg{(}1-F\big{(}r(\tau)\big{)}\bigg{)}\dot{t}(\tau)^{2}+\bigg{(}1-G\big{(}r(\tau)\big{)}\bigg{)}^{-1}\dot{r}(\tau)^{2}=-1,\qquad-\bigg{(}1-F\big{(}r(\tau)\big{)}\bigg{)}\dot{t}(\tau)=-E.$ (4)
According to the junction condition, on the gluing surface, the induced metric and the extrinsic curvature obtained from $ds_{in}^{2}$ and $ds_{out}^{2}$ respectively are equal. This leads to:
$a(\tau)^{2}\tilde{r}_{o}^{2}=r(\tau)^{2},\qquad E^{2}\frac{1-G(r(\tau))}{1-F(r(\tau))}=1.$ (5)
Combite the (5) with the (4), we can get:
$\dot{t}(\tau)=\frac{1}{E^{2}\bigg{(}1-G\big{(}r(\tau)\big{)}\bigg{)}},\qquad\dot{a}(\tau)^{2}\tilde{r}(\tau)^{2}=\dot{r}(\tau)^{2}=G\big{(}r(\tau)\big{)}.$ (6)
Moreover, the continuity of the extrinsic curvature implies $E=1$. By substituting the modified Friedmann equation (2) into the right-hand side of Eq.(6), along the same line to that in [17], a direct calculation lead us to unambiguously fix $F(r)$ and $G(r)$. The complete form of the metric (3) is now given by:
$ds_{out}^{2}=-F(r)dt^{2}+F(r)^{-1}dr^{2}+d\Omega^{2},$
where
$F(r)=\Bigg{(}1-\frac{r^{2}f(\rho)}{\gamma^{2}\Delta}\Big{(}1-f(\rho)\Big{)}(1-\frac{\rho}{\rho_{c}})\Bigg{)},\qquad f(\rho)=\frac{1-\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi r^{3}}.$ (8)
We should note that the metric (7) is valid for
$r\geq r_{b}=(\frac{3M}{4\pi\rho_{c}})^{\frac{1}{3}}.$ (9)
$r_{b}$ is the dust ball minimal radial.
In Figure 1, we investigate the effect of the mass of black hole $M$ on $F(r)$ while keeping the other parameters fixed. For simplicity, in our calculations, we set the remaining parameters as $\gamma=0.2375$. As shown in the figure, when $M$ is bigger than the extreme value $M_{b}$, the metric function $F(r)$ has two roots, corresponding to the two horizons of the black hole. When $M=M_{b}$, $F(r)$ has exactly one root, indicating the presence of a single horizon. When $M<M_{b}$, the black hole has no horizon. The critical value is found to be $M_{b}=1.2015$. For generality, in this work, we only consider the case of black holes with two horizons.
Figure 1: The metric function $F(r)$ versus r with different M.
III QUASINORMAL MODES
III.1 Massless scalar and the effective potential
To gain a deeper understanding of the spacetime properties of the metric (7), now we consider a massless scalar field perturbation in the above background. The massless scalar field $\boldsymbol{\Psi}$ is governed by the Klein-Gordon (KG) equation:
$\Box\boldsymbol{\Psi}=\frac{1}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\boldsymbol{\Psi}(t,r,\theta,\phi))=0.$ (10)
For the background metric (7) is static and spherically symmetric, we can separate the massless scalar field $\boldsymbol{\Psi}$ in the form
$\boldsymbol{\Psi}(t,r,\theta,\phi)=\sum\limits_{l,m}\frac{\Phi(t,r)}{r}Y_{l,m}(\theta,\phi),$ (11)
where $Y_{l,m}$ is the spherical harmonics, and $l$ and $m$ denote the multipole number and azimuthal number respectively. Substituting Eq. (11) into Eq. (10), the scalar equation Eq. (10) reduces to
$-\frac{\partial^{2}\Phi(t,r)}{\partial t^{2}}+F(r)^{2}\frac{\partial^{2}\Phi(t,r)}{\partial r^{2}}+F(r)F(r)^{\prime}\frac{\partial\Phi(t,r)}{\partial r}-\big{(}\frac{l(l+1)}{r^{2}}F(r)+\frac{F(r)F(r)^{\prime}}{r}\big{)}\Phi(t,r)=0.$ (12)
In order to map the radial region to (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), we define the tortoise radial coordinate $r_{*}$ by
$dr_{*}=\frac{dr}{F(r)}.$ (13)
Then Eq. (12) can be casted as
$-\frac{\partial^{2}\Phi}{\partial t^{2}}+\frac{\partial^{2}\Phi}{\partial r_{*}^{2}}-V(r)\Phi=0,$ (14)
Assuming that $\Phi(t,r)=\Psi(r)e^{-i\omega t}$, we yield
$\frac{\partial^{2}\Psi}{\partial r_{*}^{2}}+(\omega^{2}-V(r))\Psi=0.$ (15)
where the $\omega$ is the frequency of QNMs, and $V(r)$ is the effective potential which given by
$V(r)=F(r)(\frac{l(l+1)}{r^{2}}+\frac{F(r)^{\prime}}{r}).$ (16)
Physically, in the event horizon, nothing can escape, so it only have incoming waves. And the metric Eq.(7) is asymptotically flat, there are no incoming waves from infinity. Thus the boundary conditions can be chosen as
$\Psi(r)\approx e^{-i\omega r_{*}}(r\rightarrow r_{h}),\qquad\Psi(r)\approx e^{i\omega r_{*}}(r\rightarrow+\infty).$ (17)
III.2 The finite element method
In this work, we will introduce the finite element method and use it to solve the Eq. (14) to obtain the time-domain profile of $\Phi(t,r)$. We discretize the coordinates $t=i\triangle t$ and $r_{*}=j\triangle r_{*}$, where the $i$ and $j$ are integers. The scalar field $\Phi$ and the effective potential $V$ are also discretized:
$\Phi(t,r)=\Phi(i\triangle t,j\triangle r_{*})\equiv\Phi(i,j),\qquad V(r_{*})=V(j\triangle r_{*})\equiv V(j).$ (18)
So, the Eq.(14) becomes a set of iterative algebraic equations:
$-\frac{\Phi(i+1,j)-2\Phi(i,j)+\Phi(i-1,j)}{\triangle t^{2}}+\frac{\Phi(i,j+1)-2\Phi(i,j)+\Phi(i,j-1)}{\triangle r_{*}^{2}}-V(j)\Phi(i,j)=0.$ (19)
We can rewritten the Eq.(19) as
$\Phi(i+1,j)=-\Phi(i-1,j)+(2-2\frac{\triangle t^{2}}{\triangle r_{*}^{2}}-\triangle t^{2}V(j))\Phi(i,j)+\frac{\triangle t^{2}}{\triangle r_{*}^{2}}(\Phi(i,j+1)+\Phi(i,j-1)).$ (20)
The von Neumann stability condition require $\frac{\triangle t}{\triangle r_{*}}<1$, so we set $\frac{\triangle t}{\triangle r_{*}}=\frac{1}{2}$ in our numerical calculations. We consider the initial Gaussian distribution:
$\Phi(t=0,r_{*})=e^{-\frac{(r_{*}-\overline{a})^{2}}{2}},\qquad and\qquad\Phi(t<0,r_{*})=0.$ (21)
the value of $\overline{a}$ will be chosen accordingly.
III.3 WKB approximation
The WKB approximation is a semi-analytic method widely used to compute the QNMs of black holes. B. Schutz and C. Will first applied the first-order WKB method to study black hole scattering problems [47]. Subsequently, S. Iyer and C. Will extended the method to third order [48], resulting in improved accuracy, while Konoplya further pushed it to sixth order, achieving even greater precision [49]. More recently, by employing Padé approximants, Matyjasek and Opala substantially enhanced the accuracy of the WKB method and extended it up to the 13th order [50]. Then Konoplya and collaborators further improved the WKB method, significantly enhancing its precision [51].
The WKB approximation is used to solve wave-like equations where the effective potential has a barrier-like shape (a single peak) and approaches a constant as $r^{*}\to\pm\infty$. The main idea of this method is to match the approximate solutions expanded near the asymptotic regions—both near the horizon and at infinity via the two turning points, using Taylor expansions around the top of the potential barrier. Finally, for spherically symmetric backgrounds with an effective potential $V(r)$ that does not depend on the frequency $\omega$, the WKB formula can be expressed as [51]:
$\omega^{2}=$ $V_{0}+A_{2}(\mathcal{K}^{2})+A_{4}(\mathcal{K}^{2})+A_{6}(\mathcal{K}^{2})+\ldots$
$-\mathrm{i}\mathcal{K}\sqrt{-2V_{2}}(1+A_{3}(\mathcal{K}^{2})+A_{5}(\mathcal{K}^{2})+\ldots),$ (22)
where
$\mathcal{K}=n+\frac{1}{2},\quad n=0,1,2,3,\ldots$
In this expression, $n$ labels the overtone number, $V_{0}$ denotes the maximum of the effective potential $V(x)$, and $V_{n}$ refers to the $n$th derivative of $V(x)$ evaluated at this maximum. $A_{n}(\mathcal{K}^{2})$ is the $n$th-order correction, expressed as a polynomial in $\mathcal{K}^{2}$ with rational coefficients.
III.4 The QNMs results and the late-time tail
In this subsection, we investigate the angular momentum $l$ and the mass $M$ how to affect the effective potential and QNMs under scalar perturbations, and further analyze the behavior of the late-time tail. In the final stage of the scalar field evolution, the late-time tail becomes dominant. And the inverse power-law tail usually arises in the asymptotically flat spacetime.
Using the finite element method to solve equation (19), we obtain the time-domain profile. In Fig. 3 and Fig. 3, we can see that when we use the scalar waves with different angular momentum $l$ to perturb the same black hole, the height of the effective potential changes. When angular momentum $l$ increases, the peak of the effective potential increases. We should note that the tortoise radial coordinate $r_{*}$ doesn’t change the shape of the effective potential, it just stech it. The time-domain profiles correspond to this effective potential are shown in Fig. 7. As shown in the figure, we can note that the radiative decay follows the inverse power-law at late times. And when angular momentum $l$ increases, the quasinormal frequencies become longer-lived and dominate the time-domain signal.
In Fig. 5 and Fig. 5, we can see that when the mass of black hole $M$ increases, the peak of the effective potential decreases. The time-domain profiles correspond to this effective potentials shown in Figure 7. In the figure, we observe that as the mass of the LQG black hole increases, the exponential decay rate decreases and the late-time tail appears later. Furthermore, the late-time tails corresponding to different black hole masses all follow an inverse power-law behavior, indicating that the black hole mass has little effect on the late-time tail.
Fig. 9 presents the effective potential $V(r_{*})$ for the Schwarzschild black hole, the black hole in the qOS model, and the qOS model with Lorentz term respectively, all evaluated with the same mass parameter $M$ and angular momentum $l$. Near the event horizon, the effective potential exhibits slight deformations, indicating modifications in the spacetime structure of the event horizon. To further explore these structural differences, we analyze the time-domain profiles in Fig. 9. As indicated in the Figure, the profile of the black hole in the previous qOS model closely matches that of the Schwarzschild black hole, whereas the exponential decay rate in the qOS model with the Lorentz term shows a noticeable decrease. The profiles of both the Schwarzschild black hole and the black hole in the qOS model with Lorentz term display a late-time inverse power-law tail, indicating that the quantum effects of the qOS model with Lorentz term have little impact on the late-time tail behavior.
Figure 2: The effective potential $V(r)$ for different angular momentum $l$. Parameters used: $M=4.4$.
Figure 3: The effective potential $V(r_{*})$ for different angular momentum $l$. Parameters used: $M=4.4$.
Figure 4: The effective potential $V(r)$ for different mass $M$. Parameters used: $l=1$.
Figure 5: The effective potential $V(r_{*})$ for different mass $M$. Parameters used: $l=1$.
Figure 6: The time evolution of the scalar field $\Phi$ for different angular momentum $l$. Parameters used: $M=4.4$.
Figure 7: The time evolution of the scalar field $\Phi$ for different mass $M$ with $l=1$.
Figure 8: The effective potential $V(r_{*})$ for different black hole models: Schwarzschild, qOS, and qOS with Lorentz term. Parameters used: $M=1.4$, $l=1$, $\overset{\sim}{\alpha}=1.6$, where $\overset{\sim}{\alpha}$ denotes the quantum correction parameter in the qOS model.
Figure 9: The time evolution of the scalar field $\Phi$ for different black hole models: Schwarzschild, qOS, and qOS with Lorentz term. Parameters used: $M=1.4$, $l=1$, $\overset{\sim}{\alpha}=1.6$, where $\overset{\sim}{\alpha}$ denotes the quantum correction parameter in the qOS model.
III.5 Higher overtones
While the fundamental mode usually dominates the ringdown phase, it is insufficient for accurately extracting the mass and spin of the remnant black hole formed from a binary merger. This limitation can be addressed by including the first few overtones [29]. Analysis of LIGO’s observational data from the GW150914 gravitational-wave signal [52] provides evidence for the presence of the first overtone. Although corrections to Einstein’s theory must modify the geometry near the event horizon, quasinormal modes (QNMs) are generally believed to be insensitive to the near-horizon structure, as the dominant QNMs are determined primarily by the region near the peak of the effective potential. However, to better probe the geometry near the event horizon and thereby gain deeper insight into quantum gravity, it is necessary to consider the first few overtones, which are particularly sensitive to the geometric structure around the event horizon [53]. In certain scenarios, these overtones can be significantly excited and may even be detectable during the early ringdown phase by LISA [54].
To further investigate the differences in the spacetime structure among the Schwarzschild black hole, the qOS model, and the qOS model with the Lorentz term, we analyze their higher overtones using the WKB method. As shown in Table 1 and Table 2, the three models share a common trend: the real part of the frequency decreases, while the imaginary part increases with increasing overtone number $n$. Although the overall trend is consistent, significant differences in the overtone frequencies are observed for the qOS model with the Lorentz term compared to the Schwarzschild black hole, which means that $Re(\omega)$ and $Im(\omega)$ experience an outburst. Some quantum-corrected black holes, higher-derivative gravity models, and the RN black hole exhibit similar behavior [55, 56, 57]. We also find that the overtone frequencies of the qOS model with the Lorentz term show significant differences from those of the qOS model, which means that the Lorentz term has an effect on the structure of the event horizon, and this effect is reflected in the higher overtones.
To this aim, in order to gain a better understanding how the qOS model with the Lorentz term affects the spacetime structure, we analyze how the quantum-corrected parameter $\alpha$ influences the higher overtones. Here, we set $\alpha=\frac{8\gamma^{2}(1+\gamma^{2})\Delta}{M^{2}}$, which is a dimensionless parameter. The maximum value of the quantum-corrected parameter $\alpha$ is determined by the extremal value $M_{b}$. In Fig. 11 and Fig. 13, we observe that for $l=1$ and $l=2$, $Re(\omega)$ increases monotonically with increasing $\alpha$. However, in Fig. 11, the $Re(\omega)$ of $n=3$ exhibits a non-monotonic behavior. In Fig. 11 and Fig. 13, we observe that for both $l=1$ and $l=2$, $Im(\omega)$ decreases monotonically as $\alpha$ increases.
Table 1: The dominant mode and first four overtones for three different black hole models: Schwarzschild, qOS, and qOS with Lorentz term. Parameters used: $M=1.4$, $l=1$, $\overset{\sim}{\alpha}=1.6$, where $\overset{\sim}{\alpha}$ denotes the quantum correction parameter in the qOS model.
$n$ qOS with Lorentz term Schwarzschild qOS
0 0.214719--0.064490i 0.209239--0.069754i 0.212885--0.066801i
1 0.193770 --0.199489i 0.188793--0.218719i 0.193923--0.207754i
2 0.158299--0.350276i 0.163875--0.386053i 0.168338--0.363045i
3 0.119553--0.523653i 0.144981--0.562810i 0.140515--0.525821i
4 0.095652--0.716769i 0.130806--0.743082i 0.109629--0.700130i
Table 2: The dominant mode and first four overtones for three different black hole models: Schwarzschild, qOS, and qOS with Lorentz term. Parameters used: $M=1.4$, $l=2$, $\overset{\sim}{\alpha}=1.6$, where $\overset{\sim}{\alpha}$ denotes the quantum correction parameter in the qOS model.
$n$ qOS with Lorentz term Schwarzschild qOS model
0 0.354598--0.064223i 0.345460--0.069113i 0.351235--0.0663599i
1 0.340954 --0.195116i 0.331322--0.211146i 0.338155--0.202051i
2 0.315535--0.332964i 0.307538--0.363254i 0.315317--0.345657i
3 0.281461--0.481173i 0.281330--0.527130i 0.287958--0.498779i
4 0.242022--0.641936i 0.257775--0.699845i 0.259685--0.659395i
Figure 10: $Re(\omega)$ as a function of $\alpha$ for the black hole of qOS model with Lorentz term for $l=1$.
Figure 11: $Im(\omega)$ as a function of $\alpha$ for the black hole of qOS model with Lorentz term for $l=1$.
Figure 12: $Re(\omega)$ as a function of $\alpha$ for the black hole of qOS model with Lorentz term for $l=2$.
Figure 13: $Im(\omega)$ as a function of $\alpha$ for the black hole of qOS model with Lorentz term for $l=2$.
IV THERMODYNAMIC PROPERTIES
In the preceding sections, we derived the metric for a spacetime without a cosmological constant and analyzed its properties. However, our Universe is undergoing an accelerated expanding phase. This leads us to investigate the effect of the cosmological constant on spacetime, as it introduces important modifications to the structure and dynamics of spacetime. Moreover, by introducing a negative cosmological constant in black hole physics will bring us richer physics such as Hawking-Page phase transition in Schwarzschild-AdS case. Hence, the inclusion of a cosmological constant in a black hole solution have significance both in the practical and the theoretical sense. In order to investigated the thermodynamic properties of the quantum black holes in the extended formalism, one usually need to extend the effective metric to include a negative cosmological constant [58, 59]. By substituting the previous Friedmann equation Eq. (2) with the modified Friedmann equation [60] that incorporates the cosmological constant, and following the same derivation process as in the preceding sections, we can derive the metric that includes the cosmological constant, which is
$ds^{2}=-F(r)dt^{2}+F(r)^{-1}dr^{2}+d\Omega^{2},$ (23)
where
$F(r)=1-\frac{r^{2}}{\gamma^{2}\Delta}f(\rho)\Big{(}1-f(\rho)\Big{)}(1-\frac{\rho}{\rho_{c}}),\qquad f(\rho)=\frac{1\pm\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi r^{3}}+\frac{\Lambda}{8\pi G}.$
As discussed earlier, there exist two distinct branches of solutions. The $f(\rho)_{-}$ branch corresponds to a geometry that asymptotically approaches a flat spacetime, while the $f(\rho)_{+}$ branch asymptotically approaches a de Sitter spacetime with an effective cosmological constant [21, 61]. Since the bounce is expected to occur in both time directions, two scenarios are possible: a transition from the $f(\rho)_{-}$ branch to the $f(\rho)_{+}$ branch, or vice versa. However, the transition from the $f(\rho)_{+}$ branch to the $f(\rho)_{-}$ branch appears to be more physically plausible. The de Sitter–like asymptotic features of the $f(\rho)_{+}$ branch motivates further investigation into its thermodynamic properties. Therefore, we investigate the thermodynamic properties of the $f(\rho)_{+}$ branch in detail.
IV.1 Anti-de Sitter spacetime
Due to the rich and intricate thermodynamic properties exhibited by AdS spacetimes, this subsection will primarily focus on the $f(\rho)_{-}$ branch has a negative cosmological constant. In particular, we will discuss the thermodynamic properties in this context, including the Hawking temperature, heat capacity, and entropy, which are central to understanding the behavior of black holes and phase transitions in AdS backgrounds.
The Hawking temperature, which describes the thermal radiation emitted by the black hole, is related to the surface gravity at the event horizon. Specifically, the Hawking temperature can be expressed as:
$T_{H}=\frac{\mathcal{K}}{2\pi}=\frac{1}{4\pi}\frac{\partial F(r)}{\partial r}\Bigg{|}_{r=r_{h}}.$ (24)
where $\mathcal{K}$ is the surface gravity of the black hole, $r_{h}$ is the event horizon radius, which is the largest root of the Eq. (23) with the $f(\rho)_{-}$ branch. In Fig. 15, we compare the Hawking temperature for two types of black holes: the LQG black hole and the classical counterpart. As shown in the graph, the temperature continuously decreases as r increases for the classical scenario, and the temperature diverges, tending towards infinity. In contrast, the LQG black hole exhibits a different behavior. Initially, the temperature increases as r grows, reaching a maximum value. Afterward, the temperature decreases, eventually approaching the same behavior as the classical black hole at large r. In some modified gravity theories, the Hawking temperature exhibits a similar behavior [58, 62].
Figure 14: Hawking temperature of the LQG and Schwarzschild black holes as a function of even horizon $r_{h}$. Parameters used: $\Lambda=-0.001$.
Figure 15: Heat capacity of the LQG and Schwarzschild black holes as a function of even horizon $r_{h}$. Parameters used: $\Lambda=-0.001$.
In the AdS space, a large black hole is thermodynamically stable when it has a positive heat capacity, and it’s instability when it has a negative heat capacity. So, the heat capacity of a black hole provides insight into the stability of black hole. The heat capacity of black hole [63] is:
$C=\big{(}\frac{\partial M}{\partial T})=\big{(}\frac{\partial M}{\partial r_{h}})/\big{(}\frac{\partial T}{\partial r_{h}}).$ (25)
From the Eq.(23) with the $f(\rho)_{-}$ branch, we can use the event horizon $r_{h}$ to express the black hole’s mass. In Fig.15, we plot the curve of the heat capacity for $r_{h}$ and compare it with the classical scenario. As shown in the graph, we can find that the LQG black hole compared with the classical scenario has an extra phase transition. Similar phenomena can be observed in some modified gravity [58].
In AdS spacetime, the negative cosmological constant is interpreted as a positive thermodynamic pressure $P$ [64, 65], with its conjugate quantity being the thermodynamic volume $V$. Specifically, the pressure is related to the cosmological constant by: $P=\frac{-\Lambda}{8\pi}$. Thus the first law of black hole thermodynamics takes the form [66]:
$dU=TdS-PdV,$ (26)
where the $U$ is the internal energy, $T$ is the Hawking temperature, $S$ denotes the entropy. From the first law, the entropy $S$ can be derived as follows:
$S=\int\frac{dM}{T}.$ (27)
In Fig. 16, we display the behavior of the ratio $\frac{S}{r_{h}}$ as a function of the horizon radius $r_{h}$ and compare the LQG black hole with the classical scenario. As shown in the graph, we can see that the ratio $\frac{S}{r_{h}}$ for the LQG black hole exhibits a non-monotonic behavior: it initially increases with $r_{h}$, then decreases, and eventually approaches the classical result at large $r_{h}$.
In general, the entropy of a classical black hole satisfy the area law, $S=\frac{A}{4}$, where $A$ is the area of the event horizon. However, in cases involving quantum corrections and higher-dimensional gravity, the area law no longer holds. Compared to the general case, an additional logarithmic correction term appears [67, 68, 58]. In our work, the entropy also contains a logarithmic correction term, which is:
$S=\frac{A}{4}+1.506\Delta\log A.$ (28)
Figure 16: The ratio $\frac{S}{r_{h}}$ of the LQG and Schwarzschild black holes as a function of even horizon $r_{h}$. Parameters used: $\Lambda=-0.001$.
IV.2 de Sitter spacetime
The epoch of $f(\rho)_{+}$ branch does not have a cosmological constant, but its quantum geometric effects are equivalent to a positive effective cosmological constant of Planckian order [21]. The effective cosmological constant is $\Lambda=\frac{3}{(1+\gamma^{2})^{2}\Delta}$, resulting in an de-Sitter epoch. To investigate how quantum geometric effects influence black hole thermodynamics, we compare the black hole solution in the $f(\rho)_{+}$ branch with the dS Schwaschild black hole, which has the same effective cosmological constant $\Lambda$. For the $f(\rho)_{+}$ branch of the Eq. (23), we define the outer horizon as the event horizon $r_{h}$.
In Fig. 18, we investigate the Hawking temperature how to effect of the event horizon. In the dS Schwaschild black hole, the Hawking temperature increases linearly with $r_{h}$. In contrast, for the LQG black hole, the Hawking temperature first decreases with increasing $r_{h}$, and then increases linearly. Moreover, the temperature increases at different rates in the two cases, indicating that although the temperature profiles have a similar overall shape, their thermodynamic behaviors may not be identical. It should be noted that the Hawking temperature is usually positive, therefore, the temperature of the classical black hole starts from zero.
We also investigate the relationship of the heat capture with the even horizon $r_{h}$ on the Fig. 18. We can see that, compared to the Schwaschild-dS black hole, the LQG black hole has an additional phase transition, which reflects the influence of quantum effects on its thermodynamic properties. At large horizon radius $r_{h}$, the heat capacity of both the Schwaschild-dS black hole and the LQG black hole becomes negative, indicating that both of them are thermodynamically unstable.
Figure 17: Hawking temperature of the LQG and the classical Schwarzschild-dS black holes as a function of event horizon $r_{h}$.
Figure 18: Heat capacity of the LQG and the classical Schwarzschild-dS black holes as a function of event horizon $r_{h}$.
V CONCLUSIONS AND DISCUSSIONS
This study systematically explores the qOS framework within Loop Quantum Cosmology (LQC) augmented by a Lorentz term, establishing its efficacy in characterizing quantum-gravitational dynamics. The model exhibits an asymmetric quantum bounce bifurcated into two phases: the $f(\rho)_{+}$ branch governing de Sitter epoch, and the $f(\rho)_{-}$ branch describing our classical post-bounce universe. Our investigation concentrates on the phenomenologically relevant $f(\rho)_{-}$ sector.
We systematically investigate how angular momentum $l$ and the mass $M$ influence on effective potential, QNMs, and the late-time tail under scalar perturbations. We find that when angular momentum $l$ increases, the peak of the effective potential increases, and the quasinormal frequencies become longer-lived and dominate the time-domain signal. When the mass of black hole $M$ increases, the peak of the effective potential decreases. The late-time tails of black holes with different masses are similar and follow an inverse power-law behavior. To gain a better understanding of the differences in the structure of spacetime for the Schwarzschild black hole, the black hole in the qOS model, and the black hole in the qOS model with Lorentz term, we compared their effective potential and found a small deformation near the even horizon. Furthermore, we compared their time-domain profiles. The profile of the black hole in the qOS model closely matches that of the Schwarzschild black hole, while the black hole in the qOS model with Lorentz term exhibits a slower exponential decay rate. All three configurations display universal inverse power-law tails behavior across all mass configurations.
To better understand the differences in their spacetime structures near the event horizon, we further analyzed their higher overtones and found significant discrepancies in the overtone frequencies, which means the quantum gravity corrections lead to the outburst in the overtones, and this quantum gravity corrections also different from the fundamental mode. We further investigated the quantum-correction parameter $\alpha$, how it affects the higher overtones, we observe that for $l=1$ and $l=2$, $Re(\omega)$ increases monotonically with increasing $\alpha$, except for the case $n=3$, where a non-monotonic behavior is observed. Meanwhile, $Im(\omega)$ decreases monotonically with increasing $\alpha$ for both $l=1$ and $l=2$.
Note that our universe is undergoing accelerated expansion, we consider the $f(\rho)_{-}$ branch epoch having the cosmological constant. At the same time, the AdS black hole has rich thermodynamic properties, so we mainly investigate the $f(\rho)_{-}$ branch that has a negative cosmological constant. For small $r_{h}$, unlike the classical case where the Hawking temperature decreases monotonically from infinity, the temperature in the LQG black hole initially increases from zero, reaches a maximum, and then decreases. As $r_{h}$ becomes large, the behavior of the LQG black hole becomes similar to the classical scenario. For the heat capacity, we find that the LQG black hole exhibits an additional phase transition at a small $r_{h}$ compared to the classical scenario. When $r_{h}$ becomes large, the behavior of the LQG black hole coincides with that of the classical case. We also investigate the entropy of the LQG black holes. Compared to the classical case, the entropy-to-radius ratio $\frac{S}{r_{h}}$ of the LQG black hole initially increases with $r_{h}$, reaches a maximum, and then decreases, eventually converging to the classical case at large $r_{h}$. This indicates that besides the classical area law $S=\frac{A}{4}$ term, the LQG black hole also has an additional logarithmic correction term due to the quantum gravity effect.
Finally, since the $f(\rho)_{+}$ branch asymptotically approaches a de Sitter spacetime, we also investigate how quantum effects influence on it’s thermodynamic properties. We find that for the LQG black hole, the Hawking temperature first decreases with increasing horizon radius $r_{h}$, and then increases linearly, while the Hawking temperature of Schwaschild-dS black hole monotonically increases with $r_{h}$. However, the rate of increase differs between the two cases. The heat capacity of the LQG black hole exhibits an additional phase transition at small $r_{h}$ compared to the Schwaschild-dS black hole. And at large horizon radius $r_{h}$, the heat capacity of the LQG black hole becomes negative, indicating that both black holes are thermodynamically unstable.
Acknowledgements.
This work is supported by National Natural Science Foundation of China (NSFC) with Grant No. 12275087.

### v2

@@
 Recent advancements in LQG have yielded significant breakthroughs, particularly in resolving the Big Bang singularity and elucidating Hawking-Bekenstein black hole entropy [5, 6, 7].
-The application of LQG principles to homogeneous, isotropic universes has established Loop Quantum Cosmology (LQC), whose hallmark achievement is replacing the classical Big Bang singularity with a quantum bounce mechanism [8, 9, 10, 11, 12, 13, 14, 15].
+The application of LQG principles to homogeneous, isotropic universes has established Loop Quantum Cosmology (LQC), whose hallmark achievement is replacing the classical Big Bang singularity with a quantum bounce mechanism [8, 9, 10, 11, 12, 13, 14, 15, 16].
 In classical gravitational theory, the Oppenheimer-Snyder (OS) model [17] provides a foundational framework for studying gravitational collapse through pressureless, homogeneous dust dynamics governed by Friedmann equations.
 While valuable, this model inherits GR’s singularity problem.
-Recent work has addressed this limitation by applying LQC corrections to the OS framework, resulting in the quantum Oppenheimer-Snyder (qOS) model [17], which has subsequently undergone extensive theoretical investigation [18, 19, 20].
+Recent work has addressed this limitation by applying LQC corrections to the OS framework, resulting in the quantum Oppenheimer-Snyder (qOS) model [18], which has subsequently undergone extensive theoretical investigation [19, 20, 21, 22].
 Within the current qOS framework, the interior dynamics of black holes are predominantly analyzed using the simplest Loop Quantum Cosmology (LQC) approach, specifically the Ashtekar-Pawlowski-Singh (APS) model [11].
@@
 Through Thiemann’s regularization technique, a refined LQC model incorporating the Lorentzian term has been developed.
-This advancement introduces significant modifications to the Friedmann equation [25, 26] manifesting an emergent de-Sitter epoch and asymmetric bounce dynamics.
+This advancement introduces significant modifications to the Friedmann equation [27, 28, 29] manifesting an emergent de-Sitter epoch and asymmetric bounce dynamics.
 This raises a pivotal question: Would substituting the APS model with this Lorentzian-enhanced LQC preserve the qOS framework’s core conclusions?
@@
 One important physical property is that the black hole perturbation dynamics: When subjected to external perturbations, black holes exhibit a distinctive gravitational wave emission pattern that unfolds in three sequential stages.
-The initial burst phase, whose intensity and duration depend on the perturbation conditions, is followed by the quasinormal mode (QNM) ringdown - a hallmark feature where spacetime oscillations decay with characteristic complex frequencies.
+The initial burst phase, whose intensity and duration depend on the perturbation conditions, is followed by the quasinormal mode (QNM) ringdown - a hallmark feature where spacetime oscillations decay with characteristic complex frequencies [30].
 These QNM signatures, recently confirmed through groundbreaking gravitational wave detections [31, 32] have become powerful tools for probing the fundamental geometry of black holes.
@@
 In particular, the recent cosmological observations of the Universe are consistent with a de Sitter (dS) spacetime, characterized by a positive cosmological constant ($\Lambda>0$).
+In this context, the physical and thermodynamic properties of de Sitter black holes have also been extensively studied [45, 46].
 Meanwhile, the success of the AdS/CFT correspondence has highlighted the theoretical importance of anti-de Sitter (AdS) spacetime, characterized by a negative cosmological constant ($\Lambda<0$).
@@
 One of the most interesting black hole phase transitions is the Hawking-Page phase transition [50].
-So, in this work, we focus on investigating the thermodynamic properties of black holes in both de Sitter and anti-de Sitter spacetimes.
+In this work, we focus on investigating the thermodynamic properties in both de Sitter and anti-de Sitter spacetimes.
 This paper is organized as follows.
@@
 The qOS model’s spacetime comprises the pressureless dust ball and the vacuum region outside the dustball, the metric tensor inside of the dustball can be described by the FRW metric as
-$ds_{in}^{2}=-d\tau^{2}+a(\tau)^{2}(d\tilde{r}^{2}+d\Omega^{2}),$ (1)
+$ds_{in}^{2}=-d\tau^{2}+a(\tau)^{2}(d\tilde{r}^{2}+\tilde{r}^{2}d\Omega^{2}),$ (1)
 where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta\ d\phi^{2}$, and $a(\tau)$ is the scale factor of the universe.
@@
 ${\rho_{c}}=\frac{3}{32\pi G\Delta\gamma^{2}(1+\gamma^{2})}$ is the critical energy density of this model.
-In the classical limit $\frac{\rho}{\rho_{c}}\ll 1$, the (2) reduce to the standard Friedmann equation.
-When $\rho=\rho_{c}$, the quantum bounce occurs and divides the evolution of the universe into asymmetric periods.
-We should note that the modified Friedmann equation has two branches.
-When $f(\rho)=\frac{1+\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$, the corresponding modified Friedmann equation describes the pre-bounce epoch.
-Although the classical model does not contain a cosmological constant, the quantum geometric effects are equivalent to produce a positive effective cosmological constant of Planckian order, resulting in a de Sitter epoch.
-Another branch is given by $f(\rho)=\frac{1-\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$, which corresponds to the post-bounce epoch.
-This epoch is characterized by a classical universe without a cosmological constant, corresponding to the universe where we live.
-So in this work, we mainly focus on the post-bounce sector.
+We should note that the modified Friedmann equation possesses two distinct branches.
+The first branch is $f(\rho)_{+}=\frac{1+\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$, though it has no cosmological constant added at the classical evolution.
+However, its quantum geometric effects are equivalent to a positive effective cosmological constant of Planckian order [23].
+The effective cosmological constant is $\Lambda_{eff}=\frac{3}{(1+\gamma^{2})^{2}\Delta}$, resulting in an de-Sitter epoch.
+When $\rho\ll\rho_{c}$, the $f(\rho)_{+}$ branch of the modified Friedmann equation reduces to
+$H^{2}_{+}=\frac{8\pi G_{\alpha}}{3}\rho+\frac{\Lambda_{eff}}{3}.$ (3)
+where $G_{\alpha}=\frac{1-5\gamma^{2}}{\gamma^{2}+1}G$. This differs from the classical case, as it includes both a positive effective cosmological constant and a modified Newton’s constant.
+The second branch is given by $f(\rho)_{-}=\frac{1-\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})}$. When $\rho\ll\rho_{c}$, it reduces to
+$H^{2}_{-}=\frac{8\pi G}{3}\rho.$ (4)
+This branch corresponds to a classical universe in which no effective cosmological constant appears.
 Outside the dust ball’s region, we assume that the spacetime is spherically symmetric and static, so the metric reads
 $ds_{out}^{2}=-F(r)dt^{2}+G(r)^{-1}dr^{2}+d\Omega^{2}.$ (5)
+It is worthy noting that in the cosmological case, the $f(\rho)_{+}$ branch is not suitable for describing our universe, and therefore the bounce proceeds from the $f(\rho)_{+}$ branch to the $f(\rho)_{-}$ branch.
+Conversely, in the collapse case, the $f(\rho)_{-}$ branch provides a more appropriate description of stellar collapse, so the bounce proceeds from the $f(\rho)_{-}$ branch to the $f(\rho)_{+}$ branch also sounds.
+Since in both scenarios we are primarily interested in post-bounce branch, the resulting exterior solutions differ accordingly.
+A detailed explanation will be provided below.
+II.1 The vacuum exterior for the $f(\rho)_{-}$ branch
+As Eq.
+(3) shows, the emergence of the effective cosmological constant and the modified Newton’s constant is not consistent with current cosmological observations.
+Therefore, it is not suitable for describing our expanding universe.
+In the cosmological scenario, the bounce transitions from the $f(\rho)_{+}$ branch to the $f(\rho)_{-}$ branch, where the universe corresponding to the $f(\rho)_{-}$ branch is the one we inhabit.
+Accordingly, we focus on the $f(\rho)_{-}$ in the following analysis.
+Having fixed the branch of the modified Friedmann equation, we can now determine the metric outside the dust ball region.
 To determine the unknown functions $F(r)$ and $G(r)$, we need to consider the Darmois-Israel junction condition [51, 52].
-Note that the coordinates $(\tau,\tilde{r},\theta,\phi)$ describe the interior of a dust ball with a radius $0<\tilde{r}<\tilde{r}_{0}$, while the exterior is described by the coordinates $(t,r,\theta,\phi)$, with $r$ being the radial coordinate in the exterior region.
-To construct the complete spacetime, we need to glue the interior region described by the dust ball to the innermost region of a spherically symmetric and static exterior spacetime.
-In the interior region, the geodesic of a freely falling particle can be described by $\tau\mapsto(\tau,\tilde{r}_{0},\theta,\phi)$, while in the exterior region, it is described by $\tau\mapsto(t(\tau),r(\tau),\theta,\phi)$, so the single surfaces are glued by the identification $(\tau,\tilde{r}_{0},\theta,\phi)\sim(t,r,\theta,\phi)$.
-From the geodesic, we can derive:
-$-\bigg{(}1-F\big{(}r(\tau)\big{)}\bigg{)}\dot{t}(\tau)^{2}+\bigg{(}1-G\big{(}r(\tau)\big{)}\bigg{)}^{-1}\dot{r}(\tau)^{2}=-1,\qquad-\bigg{(}1-F\big{(}r(\tau)\big{)}\bigg{)}\dot{t}(\tau)=-E.$ (4)
-According to the junction condition, on the gluing surface, the induced metric and the extrinsic curvature obtained from $ds_{in}^{2}$ and $ds_{out}^{2}$ respectively are equal. This leads to:
-$a(\tau)^{2}\tilde{r}_{o}^{2}=r(\tau)^{2},\qquad E^{2}\frac{1-G(r(\tau))}{1-F(r(\tau))}=1.$ (5)
-Combite the (5) with the (4), we can get:
-$\dot{t}(\tau)=\frac{1}{E^{2}\bigg{(}1-G\big{(}r(\tau)\big{)}\bigg{)}},\qquad\dot{a}(\tau)^{2}\tilde{r}(\tau)^{2}=\dot{r}(\tau)^{2}=G\big{(}r(\tau)\big{)}.$ (6)
-Moreover, the continuity of the extrinsic curvature implies $E=1$.
-By substituting the modified Friedmann equation (2) into the right-hand side of Eq.(6), along the same line to that in [17], a direct calculation lead us to unambiguously fix $F(r)$ and $G(r)$.
-The complete form of the metric (3) is now given by:
+We first consider the radial coordinate of the surface of the dust ball, denoted by $\tilde{r_{0}}$.
+Then, the hypersurface induced from Eq.(1) is
+$ds_{in}^{2}|_{\Sigma}=-d\tau^{2}+a(\tau)^{2}\tilde{r_{0}}^{2}d\Omega^{2}.$ (6)
+We assume that every point on the surface of the dust ball shares the same proper time $\tau$, which evolves as the collapse proceeds.
+For convenience, we use $\tau$ to parameterize the hypersurface in the vacuum region outside the dust ball.
+Then, we have
+$ds_{out}^{2}|_{\Sigma}=-\bigg(F\big(r(\tau)\big)\dot{t}(\tau)^{2}-G\big(r(\tau)\big)^{-1}\dot{r}(\tau)^{2}\bigg)d\tau^{2}+r(\tau)^{2}d\Omega^{2}.$ (7)
+To construct the complete spacetime, we match the interior region of the dust ball to the exterior region of a spherically symmetric, static spacetime using the Darmois–Israel junction conditions.
+The first junction condition requires the metric to be continuous across the boundary.
+Consequently, at the boundary, the metric components satisfy
+$1=\bigg(F\big(r(\tau)\big)\dot{t}(\tau)^{2}-G\big(r(\tau)\big)^{-1}\dot{r}(\tau)^{2}\bigg),\quad\quad a(\tau)=r(\tau).$ (8)
+According to the second Darmois–Israel junction condition, the extrinsic curvature must be continuous across the boundary. In the spacetime of the interior dust ball, the non-vanishing components are
+$K^{in}_{\theta\theta}=a(\tau)\tilde{r_{0}},\quad K^{in}_{\phi\phi}=a(\tau)\tilde{r_{0}}\sin^{2}\theta.$ (9)
+For the exterior vacuum spacetime, the non-zero components of the extrinsic curvature are
+$K^{out}_{\theta\theta}=rF(r(\tau))\dot{t}\sqrt{F(r(\tau))^{-1}G(r(\tau))},\quad K^{out}_{\phi\phi}=rF(r(\tau))\dot{t}\sqrt{F(r(\tau))^{-1}G(r(\tau))}\sin^{2}\theta.$ (10)
+Since the exterior spacetime is assumed to be static, $\xi^{a}=(\frac{\partial}{\partial t})^{a}$ is a Killing vector, which allows us to define the conserved energy
+$\xi_{\mu}u^{\mu}=g^{out}_{\mu\nu}\xi^{\nu}u^{\mu}=F\big(r(\tau)\big)\dot{t}(\tau)=-E,$ (11)
+where $u^{\mu}=(\dot{t},\dot{r},0,0)$.
+Combining Eq. (9) and Eq. (10), and using Eq. (11), we obtain
+$a(\tau)\tilde{r_{0}}=rF(r(\tau))\dot{t}(\tau)\sqrt{F(r(\tau))^{-1}G(r(\tau))}=rE\sqrt{F(r(\tau))^{-1}G(r(\tau))}.$ (12)
+Using Eq. (12) together with the metric continuity condition Eq. (8), we find
+$F(r(\tau))=E^{2}G(r(\tau)),\quad G(r(\tau))=1-\dot{r}^{2}.$ (13)
+Combining Eq. (13) with the modified Friedmann equation Eq. (2), we obtain
+$G(r(\tau))=1-H^{2}r^{2}=1-\frac{r^{2}}{\gamma^{2}\Delta}f(\rho)\big(1-f(\rho)\big)\big(1-\frac{\rho}{\rho_{c}}\big).$ (14)
+For simplicity and without loss generality, we can set $E=1$, which gives $F(r)=G(r)$. The exterior metric then takes the form
 $ds_{out}^{2}$ $=-F(r)dt^{2}+F(r)^{-1}dr^{2}+d\Omega^{2},$
-where
-$F(r)=\Bigg{(}1-\frac{r^{2}f(\rho)}{\gamma^{2}\Delta}\Big{(}1-f(\rho)\Big{)}(1-\frac{\rho}{\rho_{c}})\Bigg{)},\qquad f(\rho)=\frac{1-\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi r^{3}}.$ (8)
-We should note that the metric (7) is valid for
+$F(r)$ $=1-\frac{r^{2}}{2\gamma^{2}(1+\gamma^{2})\Delta}\left(1-\dfrac{1-\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}}{2(1+\gamma^{2})}\right)\left(1-\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}\right)\left(1-\dfrac{3M}{4\pi r^{3}\rho_{c}}\right).$ (15)
+It should be noted that the metric (15) is asymptotically flat, and is valid for
 $r\geq r_{b}=(\frac{3M}{4\pi\rho_{c}})^{\frac{1}{3}}.$ (16)
@@
 Figure 1: The metric function $F(r)$ versus r with different M.
+II.2 The vacuum exterior for the $f(\rho)_{+}$ branch
+In the classical Oppenheimer–Snyder (OS) model, a homogeneous dust ball collapses inevitably to a singularity.
+In contrast, in the de Sitter OS model [53] contracts to a minimal radius and undergoes a bounce, followed by expansion.
+Studying the physical properties and formation thresholds of primordial black holes formed from gravitational collapse within the LQC framework provides a valuable probe of the quantum nature of gravity and its potential connection to dark matter [54].
+Therefore, considering gravitational collapse in this context is both important and meaningful.
+In the context of stellar collapse, the interior of the dust ball is described by the metric
+$ds_{-}^{2}=-d\tau^{2}+a(\tau)^{2}(d\tilde{r}^{2}+\tilde{r}^{2}d\Omega^{2}),$ (17)
+The dynamics of $a(\tau)$ are governed by the $f(\rho)_{-}$ branch of the modified Friedmann equation, given by
+$H_{-}^{2}=(\frac{\dot{a}}{a})^{2}=\frac{1}{2\gamma^{2}(1+\gamma^{2})\Delta}\left(1-\dfrac{1-\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}}{2(1+\gamma^{2})}\right)\left(1-\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}\right)\left(1-\dfrac{3M}{4\pi r^{3}\rho_{c}}\right).$ (18)
+The dust ball starts from a nearly static configuration at large $r$, contracts gradually, and crosses the past event horizon.
+It continues collapsing until it reaches the minimal radius $r_{b}$ where the bounce occurs.
+After the bounce, the evolution transitions to the $f(\rho)_{+}$ branch, representing the expanding phase, described by the metric
+$ds_{+}^{2}=-d\tau^{2}+a(\tau)^{2}(d\tilde{r}^{2}+\tilde{r}^{2}d\Omega^{2}),$ (19)
+with the scale factor $a(\tau)$ governed by the $f(\rho)_{+}$ branch:
+$H_{+}^{2}=(\frac{\dot{a}}{a})^{2}=\frac{1}{2\gamma^{2}(1+\gamma^{2})\Delta}\left(1-\dfrac{1+\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}}{2(1+\gamma^{2})}\right)\left(1+\sqrt{1-\dfrac{3M}{4\pi r^{3}\rho_{c}}}\right)\left(1-\dfrac{3M}{4\pi r^{3}\rho_{c}}\right).$ (20)
+The dust ball then expands outward, crosses the future cosmological horizon, and continues to infinity.
+It should be noted that only the $f(\rho)_{-}$ branch reproduces the classical Oppenheimer–Snyder collapse in the $\rho\ll\rho_{c}$, while the $f(\rho)_{+}$ branch does not, as discussed previously.
+Consequently, the big bounce corresponds to a transition from the $f(\rho)_{-}$ branch to the $f(\rho)_{+}$ branch, and we are primarily interested in post-bounce branch.
+Once the appropriate branch of the modified Friedmann equation is identified, the exterior metric of the dust ball can be determined following the same procedure as outlined before, resulting in the metric
+$ds_{out}^{2}$ $=-F(r)dt^{2}+F(r)^{-1}dr^{2}+d\Omega^{2},$
+$F(r)$ $=1-\frac{r^{2}}{2\gamma^{2}(1+\gamma^{2})\Delta}\left(1-\frac{1+\sqrt{1-\frac{3M}{4\pi r^{3}\rho_{c}}}}{2(1+\gamma^{2})}\right)\left(1+\sqrt{1-\frac{3M}{4\pi r^{3}\rho_{c}}}\right)\left(1-\frac{3M}{4\pi r^{3}\rho_{c}}\right).$ (21)
 III QUASINORMAL MODES
+As mentioned earlier, the vacuum exterior of the $f(\rho)_{-}$ branch corresponds to a spacetime containing a black hole, whereas the $f(\rho)_{+}$ branch admits only a cosmological horizon.
+Moreover, the vacuum exterior of the $f(\rho)_{-}$ branch appears to be more consistent with cosmological observations.
+Therefore, in this section, we restrict our analysis to the $f(\rho)_{-}$ branch.
 III.1 Massless scalar and the effective potential
@@
 $-\frac{\partial^{2}\Phi(t,r)}{\partial t^{2}}+F(r)^{2}\frac{\partial^{2}\Phi(t,r)}{\partial r^{2}}+F(r)F(r)^{\prime}\frac{\partial\Phi(t,r)}{\partial r}-\big(\frac{l(l+1)}{r^{2}}F(r)+\frac{F(r)F(r)^{\prime}}{r}\big)\Phi(t,r)=0.$ (24)
-In order to map the radial region to (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), we define the tortoise radial coordinate $r_{*}$ by
+In order to map the radial region to $(-\infty,+\infty)$, we define the tortoise radial coordinate $r_{*}$ by
 $dr_{*}=\frac{dr}{F(r)}.$ (25)
@@
 Physically, in the event horizon, nothing can escape, so it only have incoming waves.
-And the metric Eq.(7) is asymptotically flat, there are no incoming waves from infinity.
+And the metric Eq.(LABEL:eq:_new_metric) is asymptotically flat, there are no incoming waves from infinity.
 Thus the boundary conditions can be chosen as
@@
 The time-domain profiles correspond to this effective potentials shown in Figure 7.
-In the figure, we observe that as the mass of the LQG black hole increases, the exponential decay rate decreases and the late-time tail appears later.
+In the figure, we observe that as the mass of the LQG-inspired black hole increases, the exponential decay rate decreases and the late-time tail appears later.
 Furthermore, the late-time tails corresponding to different black hole masses all follow an inverse power-law behavior, indicating that the black hole mass has little effect on the late-time tail.
@@
 As indicated in the Figure, the profile of the black hole in the previous qOS model closely matches that of the Schwarzschild black hole, whereas the exponential decay rate in the qOS model with the Lorentz term shows a noticeable decrease.
+It should be noted that the deformation induced by quantum effects is minimal compared to the classical case.
 The profiles of both the Schwarzschild black hole and the black hole in the qOS model with Lorentz term display a late-time inverse power-law tail, indicating that the quantum effects of the qOS model with Lorentz term have little impact on the late-time tail behavior.
@@
 $n$ qOS with Lorentz term Schwarzschild qOS
-0 0.214719--0.064490i 0.209239--0.069754i 0.212885--0.066801i
-1 0.193770 --0.199489i 0.188793--0.218719i 0.193923--0.207754i
-2 0.158299--0.350276i 0.163875--0.386053i 0.168338--0.363045i
-3 0.119553--0.523653i 0.144981--0.562810i 0.140515--0.525821i
-4 0.095652--0.716769i 0.130806--0.743082i 0.109629--0.700130i
+0 0.214719-0.064490i 0.209239-0.069754i 0.212885-0.066801i
+1 0.193770 -0.199489i 0.188793-0.218719i 0.193923-0.207754i
+2 0.158299-0.350276i 0.163875-0.386053i 0.168338-0.363045i
+3 0.119553-0.523653i 0.144981-0.562810i 0.140515-0.525821i
+4 0.095652-0.716769i 0.130806-0.743082i 0.109629-0.700130i
 Table 2: The dominant mode and first four overtones for three different black hole models: Schwarzschild, qOS, and qOS with Lorentz term.
@@
 $n$ qOS with Lorentz term Schwarzschild qOS model
-0 0.354598--0.064223i 0.345460--0.069113i 0.351235--0.0663599i
-1 0.340954 --0.195116i 0.331322--0.211146i 0.338155--0.202051i
-2 0.315535--0.332964i 0.307538--0.363254i 0.315317--0.345657i
-3 0.281461--0.481173i 0.281330--0.527130i 0.287958--0.498779i
-4 0.242022--0.641936i 0.257775--0.699845i 0.259685--0.659395i
+0 0.354598-0.064223i 0.345460-0.069113i 0.351235-0.0663599i
+1 0.340954 -0.195116i 0.331322-0.211146i 0.338155-0.202051i
+2 0.315535-0.332964i 0.307538-0.363254i 0.315317-0.345657i
+3 0.281461-0.481173i 0.281330-0.527130i 0.287958-0.498779i
+4 0.242022-0.641936i 0.257775-0.699845i 0.259685-0.659395i
 Figure 10: $Re(\omega)$ as a function of $\alpha$ for the black hole of qOS model with Lorentz term for $l=1$.
@@
 IV THERMODYNAMIC PROPERTIES
-In the preceding sections, we derived the metric for a spacetime without a cosmological constant and analyzed its properties.
-However, our Universe is undergoing an accelerated expanding phase.
-This leads us to investigate the effect of the cosmological constant on spacetime, as it introduces important modifications to the structure and dynamics of spacetime.
-Moreover, by introducing a negative cosmological constant in black hole physics will bring us richer physics such as Hawking-Page phase transition in Schwarzschild-AdS case.
-Hence, the inclusion of a cosmological constant in a black hole solution have significance both in the practical and the theoretical sense.
-In order to investigated the thermodynamic properties of the quantum black holes in the extended formalism, one usually need to extend the effective metric to include a negative cosmological constant [58, 59].
+In the preceding sections, we derived the metric corresponding to the post-bounce branch of the spacetime without a cosmological constant for both the cosmological and collapse scenarios, and analyzed its properties.
+As discussed earlier, the metric of the spacetime on the post-bounce branch differs between the cosmological and collapse cases.
+In the cosmological case, the vacuum exterior metric corresponding to the post-bounce branch is denoted by $f(\rho)_{-}$, whose geometry asymptotically approaches flat spacetime.
+In contrast, for the collapse case, the vacuum exterior metric corresponding to the post-bounce branch is given by $f(\rho)_{+}$, which asymptotically approaches a de Sitter spacetime with an effective cosmological constant [23, 29].
+However, our Universe is currently undergoing an accelerated expansion, the simplest possible explanation of which is the presence of a positive cosmological constant ($\Lambda>0$).
+This motivates us to investigate the effect of the cosmological constant on the spacetime, as it introduces important modifications to its structure and dynamics.
+Moreover, the success of the AdS/CFT correspondence has drawn considerable attention to the case with a negative cosmological constant ($\Lambda<0$).
+Introducing a negative cosmological constant in black hole physics also leads to richer phenomena.
+Hence, the inclusion of a cosmological constant in the solution have significance both in the practical and the theoretical sense.
+To investigate the thermodynamic properties of the vacuum exterior metrics for the $f(\rho)_{-}$ and $f(\rho)_{+}$ branches, it is necessary to extend the metrics to include the cosmological constant term [66, 67].
 By substituting the previous Friedmann equation Eq.
@@
 where
-$F(r)=1-\frac{r^{2}}{\gamma^{2}\Delta}f(\rho)\Big{(}1-f(\rho)\Big{)}(1-\frac{\rho}{\rho_{c}}),\qquad f(\rho)=\frac{1\pm\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi r^{3}}+\frac{\Lambda}{8\pi G}.$
-As discussed earlier, there exist two distinct branches of solutions.
-The $f(\rho)_{-}$ branch corresponds to a geometry that asymptotically approaches a flat spacetime, while the $f(\rho)_{+}$ branch asymptotically approaches a de Sitter spacetime with an effective cosmological constant [21, 61].
-Since the bounce is expected to occur in both time directions, two scenarios are possible: a transition from the $f(\rho)_{-}$ branch to the $f(\rho)_{+}$ branch, or vice versa.
-However, the transition from the $f(\rho)_{+}$ branch to the $f(\rho)_{-}$ branch appears to be more physically plausible.
-The de Sitter–like asymptotic features of the $f(\rho)_{+}$ branch motivates further investigation into its thermodynamic properties.
-Therefore, we investigate the thermodynamic properties of the $f(\rho)_{+}$ branch in detail.
+$F(r)=1-\frac{r^{2}}{\gamma^{2}\Delta}f(\rho)\Big(1-f(\rho)\Big)(1-\frac{\rho}{\rho_{c}}),\qquad f(\rho)_{\pm}=\frac{1\pm\sqrt{1-\frac{\rho}{\rho_{c}}}}{2(1+\gamma^{2})},\qquad\rho=\frac{M}{\frac{4}{3}\pi r^{3}}+\frac{\Lambda}{8\pi G}.$
+To capture the richer thermodynamic features of the black hole, we incorporate a negative cosmological constant into the vacuum exterior metric of the $f(\rho)_{-}$ branch.
+As discussed earlier, the metric of the $f(\rho)_{+}$ branch does not contain a cosmological constant, but its quantum geometric effects are equivalent to a positive effective cosmological constant, and resulting in an de-Sitter epoch.
+To examine how the quantum geometric effects influence the thermodynamic properties, we compare the thermodynamic behavior of this spacetime with that of the classical case.
+Furthermore, as we show below, incorporating a negative cosmological constant into the metric of this case does not alter our main conclusions.
 IV.1 Anti-de Sitter spacetime
-Due to the rich and intricate thermodynamic properties exhibited by AdS spacetimes, this subsection will primarily focus on the $f(\rho)_{-}$ branch has a negative cosmological constant.
+Due to the rich and intricate thermodynamic properties exhibited by AdS spacetimes, this subsection will primarily focus on the metric of the $f(\rho)_{-}$ branch with a negative cosmological constant.
 In particular, we will discuss the thermodynamic properties in this context, including the Hawking temperature, heat capacity, and entropy, which are central to understanding the behavior of black holes and phase transitions in AdS backgrounds.
@@
 In Fig.
-15, we compare the Hawking temperature for two types of black holes: the LQG black hole and the classical counterpart.
-As shown in the graph, the temperature continuously decreases as r increases for the classical scenario, and the temperature diverges, tending towards infinity.
-In contrast, the LQG black hole exhibits a different behavior.
-Initially, the temperature increases as r grows, reaching a maximum value.
-Afterward, the temperature decreases, eventually approaching the same behavior as the classical black hole at large r.
+15, we compare the Hawking temperature for two types of black holes: the LQG-inspired black hole and the classical counterpart.
+As shown in the graph, the temperature monotonically decreases as $r_{h}$ increases in the classical scenario.
+When $r_{h}$ approaches zero, the temperature diverges, tending towards infinity.
+In contrast, the LQG-inspired black holes exhibits a different behavior.
+The temperature starts at the extremal case, where the $f(\rho)_{-}$ branch of the metric in Eq.
+(35) has a single horizon.
+Initially, the temperature increases as $r_{h}$ grows, reaching a maximum value.
+Afterward, the temperature decreases, eventually approaching the same behavior as the classical black hole at large $r_{h}$.
 In some modified gravity theories, the Hawking temperature exhibits a similar behavior [66, 69].
@@
 In Fig.15, we plot the curve of the heat capacity for $r_{h}$ and compare it with the classical scenario.
-As shown in the graph, we can find that the LQG black hole compared with the classical scenario has an extra phase transition.
+As shown in the graph, we can find that the LQG-inspired black hole compared with the classical scenario has an extra phase transition.
 Similar phenomena can be observed in some modified gravity [66].
@@
 In Fig.
-16, we display the behavior of the ratio $\frac{S}{r_{h}}$ as a function of the horizon radius $r_{h}$ and compare the LQG black hole with the classical scenario.
-As shown in the graph, we can see that the ratio $\frac{S}{r_{h}}$ for the LQG black hole exhibits a non-monotonic behavior: it initially increases with $r_{h}$, then decreases, and eventually approaches the classical result at large $r_{h}$.
-In general, the entropy of a classical black hole satisfy the area law, $S=\frac{A}{4}$, where $A$ is the area of the event horizon.
+16, we display the behavior of the ratio $\frac{S}{A}$ as a function of the horizon radius $r_{h}$ and compare the LQG-inspired black hole with the classical scenario, where $A$ is the area of the event horizon.
+As shown in the graph, we can see that the ratio $\frac{S}{A}$ for the LQG-inspired black hole exhibits a non-monotonic behavior: it initially increases with $r_{h}$, then decreases, and eventually approaches the classical result at large $r_{h}$.
+In general, the entropy of a classical black hole satisfy the area law, $S=\frac{A}{4}$.
 However, in cases involving quantum corrections and higher-dimensional gravity, the area law no longer holds.
@@
 $S=\frac{A}{4}+1.506\Delta\log A.$ (40)
-Figure 16: The ratio $\frac{S}{r_{h}}$ of the LQG and Schwarzschild black holes as a function of even horizon $r_{h}$. Parameters used: $\Lambda=-0.001$.
+Figure 16: The ratio $\frac{S}{A}$ of the LQG and Schwarzschild black holes as a function of even horizon $r_{h}$. Parameters used: $\Lambda=-0.001$.
 IV.2 de Sitter spacetime
-The epoch of $f(\rho)_{+}$ branch does not have a cosmological constant, but its quantum geometric effects are equivalent to a positive effective cosmological constant of Planckian order [21].
-The effective cosmological constant is $\Lambda=\frac{3}{(1+\gamma^{2})^{2}\Delta}$, resulting in an de-Sitter epoch.
-To investigate how quantum geometric effects influence black hole thermodynamics, we compare the black hole solution in the $f(\rho)_{+}$ branch with the dS Schwaschild black hole, which has the same effective cosmological constant $\Lambda$.
+The metric of the $f(\rho)_{+}$ branch does not contain a cosmological constant, but its quantum geometric effects are equivalent to a positive effective cosmological constant of Planckian order [23].
+The effective cosmological constant is $\Lambda_{eff}=\frac{3}{(1+\gamma^{2})^{2}\Delta}$, resulting in an de-Sitter epoch.
+To investigate how quantum geometric effects influence its thermodynamics, we compare this solution with the de Sitter Schwarzschild black hole, which has the same effective cosmological constant $\Lambda_{eff}$.
 For the $f(\rho)_{+}$ branch of the Eq.
-(23), we define the outer horizon as the event horizon $r_{h}$.
+(35), we define the cosmological horizon as $r_{c}$.
 In Fig.
-18, we investigate the Hawking temperature how to effect of the event horizon.
-In the dS Schwaschild black hole, the Hawking temperature increases linearly with $r_{h}$.
-In contrast, for the LQG black hole, the Hawking temperature first decreases with increasing $r_{h}$, and then increases linearly.
+18, we investigate the Hawking temperature how to effect of the cosmological horizon.
+In the de Sitter Schwarzschild black hole, the Hawking temperature increases linearly with $r_{c}$.
+In contrast, for the LQG-inspired black hole, the Hawking temperature first decreases with increasing $r_{c}$, and then increases linearly.
 Moreover, the temperature increases at different rates in the two cases, indicating that although the temperature profiles have a similar overall shape, their thermodynamic behaviors may not be identical.
 It should be noted that the Hawking temperature is usually positive, therefore, the temperature of the classical black hole starts from zero.
-We also investigate the relationship of the heat capture with the even horizon $r_{h}$ on the Fig.
+We also investigate the relationship of the heat capture with the cosmological horizon $r_{c}$ on the Fig.
 18.
-We can see that, compared to the Schwaschild-dS black hole, the LQG black hole has an additional phase transition, which reflects the influence of quantum effects on its thermodynamic properties.
-At large horizon radius $r_{h}$, the heat capacity of both the Schwaschild-dS black hole and the LQG black hole becomes negative, indicating that both of them are thermodynamically unstable.
-Figure 17: Hawking temperature of the LQG and the classical Schwarzschild-dS black holes as a function of event horizon $r_{h}$.
-Figure 18: Heat capacity of the LQG and the classical Schwarzschild-dS black holes as a function of event horizon $r_{h}$.
+We can see that, compared to the de Sitter Schwarzschild black hole, the LQG solution has an additional phase transition, which reflects the influence of quantum effects on its thermodynamic properties.
+At large horizon radius $r_{c}$, the heat capacity of both the de Sitter Schwarzschild black hole and the LQG solution becomes negative, indicating that both of them are thermodynamically unstable.
+Figure 17: Hawking temperature of the LQG and the classical Schwarzschild-dS black holes as a function of cosmological horizon $r_{c}$.
+Figure 18: Heat capacity of the LQG and the classical Schwarzschild-dS black holes as a function of cosmological horizon $r_{c}$.
+We can also incorporate a negative cosmological constant into this metric, while ensuring that its absolute value does not exceed the effective cosmological constant $\Lambda_{eff}$.
+We find that the conclusion remains unchanged.
 V CONCLUSIONS AND DISCUSSIONS
 This study systematically explores the qOS framework within Loop Quantum Cosmology (LQC) augmented by a Lorentz term, establishing its efficacy in characterizing quantum-gravitational dynamics.
-The model exhibits an asymmetric quantum bounce bifurcated into two phases: the $f(\rho)_{+}$ branch governing de Sitter epoch, and the $f(\rho)_{-}$ branch describing our classical post-bounce universe.
-Our investigation concentrates on the phenomenologically relevant $f(\rho)_{-}$ sector.
+The model exhibits an asymmetric quantum bounce bifurcated into two phases: the $f(\rho)_{+}$ branch governing de Sitter epoch, and the $f(\rho)_{-}$ branch asymptotic flat.
+In the cosmological case, the bounce allow to transit from $f(\rho)_{+}$ branch to $f(\rho)_{-}$ branch, whereas in the collapse scenario,it transitions from $f(\rho)_{-}$ branch to $f(\rho)_{+}$ branch.
+In both cases, the post-bounce branch describes the universe in which we currently live.
+Consequently, our investigation focuses on the phenomenologically relevant post-bounce sector.
+In the cosmological case, the vacuum exterior associated with the post-bounce $f(\rho)_{-}$ branch contains a black hole, whereas in the collapse scenario, the vacuum exterior corresponding to the post-bounce $f(\rho)_{+}$ branch features only a cosmological horizon.
+Therefore, in what follows, we focus our quasinormal-mode (QNM) analysis on the $f(\rho)_{-}$ branch.
 We systematically investigate how angular momentum $l$ and the mass $M$ influence on effective potential, QNMs, and the late-time tail under scalar perturbations.
@@
 Meanwhile, $Im(\omega)$ decreases monotonically with increasing $\alpha$ for both $l=1$ and $l=2$.
-Note that our universe is undergoing accelerated expansion, we consider the $f(\rho)_{-}$ branch epoch having the cosmological constant.
-At the same time, the AdS black hole has rich thermodynamic properties, so we mainly investigate the $f(\rho)_{-}$ branch that has a negative cosmological constant.
-For small $r_{h}$, unlike the classical case where the Hawking temperature decreases monotonically from infinity, the temperature in the LQG black hole initially increases from zero, reaches a maximum, and then decreases.
-As $r_{h}$ becomes large, the behavior of the LQG black hole becomes similar to the classical scenario.
-For the heat capacity, we find that the LQG black hole exhibits an additional phase transition at a small $r_{h}$ compared to the classical scenario.
-When $r_{h}$ becomes large, the behavior of the LQG black hole coincides with that of the classical case.
-We also investigate the entropy of the LQG black holes.
-Compared to the classical case, the entropy-to-radius ratio $\frac{S}{r_{h}}$ of the LQG black hole initially increases with $r_{h}$, reaches a maximum, and then decreases, eventually converging to the classical case at large $r_{h}$.
-This indicates that besides the classical area law $S=\frac{A}{4}$ term, the LQG black hole also has an additional logarithmic correction term due to the quantum gravity effect.
-Finally, since the $f(\rho)_{+}$ branch asymptotically approaches a de Sitter spacetime, we also investigate how quantum effects influence on it’s thermodynamic properties.
-We find that for the LQG black hole, the Hawking temperature first decreases with increasing horizon radius $r_{h}$, and then increases linearly, while the Hawking temperature of Schwaschild-dS black hole monotonically increases with $r_{h}$.
+The cosmological constant is generally believed to be the simplest explanation for the accelerated expansion of our Universe, while a negative cosmological constant introduces rich and distinctive physical signatures.
+Therefore, we consider the vacuum exterior metrics with a negative cosmological constant for both the $f(\rho)_{-}$ and $f(\rho)_{+}$ branches, and investigate their thermodynamic properties.
+For the vacuum exterior metric corresponding to the $f(\rho)_{-}$ branch with a negative cosmological constant, we find that for small $r_{h}$, unlike the classical case where the Hawking temperature decreases monotonically from infinity, the temperature in the LQG-inspired black hole initially increases from zero, reaches a maximum, and then decreases.
+As $r_{h}$ becomes large, the behavior of the LQG-inspired black hole becomes similar to the classical scenario.
+For the heat capacity, we find that the LQG-inspired black hole exhibits an additional phase transition at a small $r_{h}$ compared to the classical scenario.
+When $r_{h}$ becomes large, the behavior of the LQG-inspired black hole coincides with that of the classical case.
+We also investigate the entropy of the LQG-inspired black hole.
+Compared to the classical case, the entropy-to-area ratio $\frac{S}{A}$ of the LQG-inspired black hole initially increases with $r_{h}$, reaches a maximum, and then decreases, eventually converging to the classical case at large $r_{h}$.
+This indicates that besides the classical area law $S=\frac{A}{4}$ term, the LQG-inspired black hole also has an additional logarithmic correction term due to the quantum gravity effect.
+For the vacuum exterior metric corresponding to the $f(\rho)_{+}$ branch, we first consider the spacetime without a cosmological constant and compare its thermodynamic behavior with that of the classical Schwarzschild case.
+Since the quantum geometric effects in this model are equivalent to a positive effective cosmological constant $\Lambda_{eff}$, this comparison allows us to examine how such effects influence the thermodynamic properties.
+We find that for the LQG-inspired black hole, the Hawking temperature first decreases with increasing horizon radius $r_{c}$, and then increases linearly, while the Hawking temperature of de Sitter Schwarzschild black hole monotonically increases with $r_{c}$.
 However, the rate of increase differs between the two cases.
-The heat capacity of the LQG black hole exhibits an additional phase transition at small $r_{h}$ compared to the Schwaschild-dS black hole.
-And at large horizon radius $r_{h}$, the heat capacity of the LQG black hole becomes negative, indicating that both black holes are thermodynamically unstable.
+The heat capacity of the LQG-inspired black hole exhibits an additional phase transition at small $r_{h}$ compared to the de Sitter Schwarzschild black hole.
+And at large horizon radius $r_{c}$, the heat capacity of the LQG-inspired black hole becomes negative, indicating that both black holes are thermodynamically unstable.
+Furthermore, as discussed above, even when a negative cosmological constant is introduced into the metric, provided that its absolute value does not exceed the effective cosmological constant $\Lambda_{eff}$, our main conclusions remain unchanged.
 Acknowledgements.